OpenStudy (anonymous):

Write the inequality for a graph that has the boundary line is dashed, a slope of 1/3, the y-intercept -2, and the area above the line is shaded.

OpenStudy (anonymous):

Can you help @RadEn??

OpenStudy (raden):

ok, first we assume it is a straight line equation y=mx+c according info above, m=1/3 and c=-2. so, the line is y=1/3*x - 2 right ?

OpenStudy (anonymous):

right

OpenStudy (raden):

ok, then we can convert to inequality sign (it can be > or <) because the area above the line is shaded, therefore gave sign ">" ao, y>(1/3)x - 2

OpenStudy (raden):

*so

OpenStudy (anonymous):

thank you!

OpenStudy (raden):

welcome :)